OpenStudy (anonymous):

Marnie packed 18 boxes in 3 hours, which was 72% of the total number of boxes she had to pack. How many boxes did Marnie have to pack?

OpenStudy (anonymous):

id like to know how to get the answer as well!:)

OpenStudy (anonymous):

18 = (0.72) (total) total = ?

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

got it! thx. :)

OpenStudy (anonymous):

18 boxes is 72% x boxes is 100% Now use the "X" way (I don't know how you call it in English we just multiply these like a X) So 18*100=72*x 1800=72x 1800:72=x x=25 boxes

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

np